OpenStudy (jenniferjuice):

What is the vertex of the graph of y = 2(x − 3)2 + 4? (4, 3) (4, −3) (3, 4) (−3, 4)

OpenStudy (anonymous):

Recall that the vertex form of a parabola is \(\displaystyle y=a(x-h)^2+k\) where \((h,k)\) is the vertex. So in your case, what is \((h,k)\)? :-)

OpenStudy (aum):

Compare: \(\large y = a(x-h)^2 + k\) and \(\large y = 2(x-3)^2 + 4\) h = ? k = ?

OpenStudy (jenniferjuice):

h=3 k=4

OpenStudy (anonymous):

Correct, so the vertex is (3,4). I hope that made sense!